OpenStudy (anonymous):

-√-4 Is this just 2?

jimthompson5910 (jim_thompson5910):

the square root of negative number is not a real number

OpenStudy (anonymous):

jim_thompson5910 thank you

jimthompson5910 (jim_thompson5910):

sqrt(-4) = sqrt(-1*4) sqrt(-4) = sqrt(-1)*sqrt(4) sqrt(-4) = i*2 sqrt(-4) = 2i

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

how about -√500? I got -10√5

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

awesome!

jimthompson5910 (jim_thompson5910):

oh sry, made a typo should be -sqrt(-4) = -2i

jimthompson5910 (jim_thompson5910):

but I'm sure you figured that out

OpenStudy (anonymous):

√81x^6 I got 9√(x)^3

jimthompson5910 (jim_thompson5910):

no just 9x^3

OpenStudy (anonymous):

really? hmmm

jimthompson5910 (jim_thompson5910):

when you take the square root of x^6, you get x^3 and the root goes away

OpenStudy (anonymous):

oooooooh. Alright I get it

OpenStudy (anonymous):

√64x^16y^34

OpenStudy (anonymous):

I got4√2(x^2)(y^17)

jimthompson5910 (jim_thompson5910):

square root of 64 = ??

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

square root of x^16 = ??

OpenStudy (anonymous):

then I went further with 4√2

OpenStudy (anonymous):

x^16

OpenStudy (anonymous):

is

OpenStudy (anonymous):

x^4

jimthompson5910 (jim_thompson5910):

no once you get 8, there are no more roots over 8

jimthompson5910 (jim_thompson5910):

so it's just 8

jimthompson5910 (jim_thompson5910):

for that part

OpenStudy (anonymous):

8^4

OpenStudy (anonymous):

8x^4

OpenStudy (anonymous):

8(x^2)(y^17)?

jimthompson5910 (jim_thompson5910):

no square root of x^16 is x^8

jimthompson5910 (jim_thompson5910):

square root of y^34 is y^17

jimthompson5910 (jim_thompson5910):

so the answer is 8x^8y^17

OpenStudy (anonymous):

Is it to be written that way as well?

jimthompson5910 (jim_thompson5910):

yeah pretty much

OpenStudy (anonymous):

Thank you, just practicing here

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Can you assist with this one?

OpenStudy (anonymous):

(x+5)^2=-15 solve by using square root property

OpenStudy (anonymous):

Im stuck because there is a negative and you change a positive into a negative so I stuck

jimthompson5910 (jim_thompson5910):

(x+5)^2=-15 x+5 = sqrt(-15) or x+5 = -sqrt(-15) x = 5+sqrt(-15) or x = -5-sqrt(-15) x = 5+i*sqrt(15) or x = -5-i*sqrt(15)

OpenStudy (anonymous):

So its the same thing in the beginning?

OpenStudy (anonymous):

x+5 = sqrt(-15) or x+5 = -sqrt(-15)

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

Ooooooooooh I just saw the difference! -sqrt(-15) and sqrt(-15)

jimthompson5910 (jim_thompson5910):

yeah it's hard to spot sometimes

OpenStudy (anonymous):

I have not learned anything with the i's yet so I put no solution

jimthompson5910 (jim_thompson5910):

ok that works

OpenStudy (anonymous):

thanks again1

jimthompson5910 (jim_thompson5910):

np